1)     If the total production function given as: TP = 16L2 - 0.4L3 then;

A)   Find the maximum value of APL and MPL.

Calculate the maximum production 


1
Expert's answer
2022-01-12T09:10:45-0500

Given "TP=16L^2-0.4L^3", we find the maximum value of APL and MPL as follows.

"1)"

"APL = \frac{TP}{L}"

"APL = \frac{( 16L^{2}-0.4L^{3})}{L}"

"APL = 16L -0.4L^{2}"

The maximum value of APL can be determined by setting "APL=0"

Therefore,

"16L -0.4L^{2}=0"

"16L = 0.4L^2\implies16=0.4L"

"L = 40"

The maximum value of APL is 40.


"2)"

The MPL is found by differentiating TP with respect to L.

"MPL=TP'=32L-1.2L^2".

The maximum value of MPL is found by setting MPL=0

So,

"32L-1.2L^2=0\implies32=1.2L\implies L=26.6667"

Therefore, the maximum value of APL and MPL will be 40 and 26.6667 respectively.


"3)"

The Maximum production can be determined by substituting the maximum value of MPL of 26.6667 in the total production function as follows.

"TP = 16L^{2}-0.4L^{3}"

"TP = 16\times(26.6667)^2 - 0.4\times(26.6667)^3"

"TP = 3792.6"

The Maximum production will be 3792.6